# 2Huawei Noah’s Ark Lab, London, UK

### v1

@@
11institutetext: 🖂 Corresponding authors
1School of Informatics, University of Edinburgh, Edinburgh, UK
2Huawei Noah’s Ark Lab, London, UK
Learning Precise Affordances from Egocentric Videos for Robotic Manipulation
Gen Li1    Nikolaos Tsagkas1    Jifei Song2    Ruaridh Mon-Williams1    Sethu Vijayakumar1    Kun Shao2 🖂    Laura Sevilla-Lara1 🖂
(Received: date / Accepted: date)
Zusammenfassung
Affordance, defined as the potential actions that an object offers, is crucial for robotic manipulation tasks. A deep understanding of affordance can lead to more intelligent AI systems. For example, such knowledge directs an agent to grasp a knife by the handle for cutting and by the blade when passing it to someone. In this paper, we present a streamlined affordance learning system that encompasses data collection, effective model training, and robot deployment. First, we collect training data from egocentric videos in an automatic manner. Different from previous methods that focus only on the object graspable affordance and represent it as coarse heatmaps, we cover both graspable (e. g., object handles) and functional affordances (e. g., knife blades, hammer heads) and extract data with precise segmentation masks. We then propose an effective model, termed Geometry-guided Affordance Transformer (GAT), to train on the collected data. GAT integrates an innovative Depth Feature Injector (DFI) to incorporate 3D shape and geometric priors, enhancing the model’s understanding of affordances. To enable affordance-oriented manipulation, we further introduce Aff-Grasp, a framework that combines GAT with a grasp generation model. For comprehensive evaluation, we create an affordance evaluation dataset with pixel-wise annotations, and design real-world tasks for robot experiments. The results show that GAT surpasses the state-of-the-art by 15.9% in mIoU, and Aff-Grasp achieves high success rates of 95.5% in affordance prediction and 77.1% in successful grasping among 179 trials, including evaluations with seen, unseen objects, and cluttered scenes. Project page: https://reagan1311.github.io/affgrasp.
Keywords:
Affordance learning Egocentric videos Human-object interaction Robotic manipulation
journal: International Journal of Computer Vision
@@ 1 Introduction
Humans excel at mastering the use of various tools and selecting the appropriate one based on its function and utility. At a dining table, for instance, we use a knife to cut, a fork to stick, and a spoon to scoop. In addition to recognizing the appearance and semantics of tools, we know how to grasp them and which part to use for the desired function. This capability is driven by affordance, which denotes the potential actions that objects offer gibson . With this understanding, we can effectively manipulate and use a variety of objects, even in unfamiliar environments. Similarly, equipping robots with this actionable knowledge is crucial for intelligent interactions, making it a popular research topic in the fields of computer vision and robotics.
A broad spectrum of work 3d-affordancenet ; nguyendetecting ; affordancenet ; learn2act_properly ; aff_with_CNN_iit ; bay-aff ; tool_parts_umd focuses on learning from hand-crafted affordance datasets, which require extensive data collection and costly annotation. In contrast, humans typically acquire knowledge of affordances in a more efficient manner, either through trial-and-error interactions or by observing others. Intuitively, learning through observation is particularly effective, allowing for generalization to objects with similar shapes and appearances. Inspired by this intuition, a number of recent studies demo2vec ; grounded ; joint_hand_hotspot ; vrb ; robo-abc ; text-aff have focused on extracting actionable knowledge from videos of humans interacting with objects. These studies reason about object affordances from human videos and represent them as interaction heatmaps. However, two limitations persist in their learning pipeline as illustrated in Figure 1: (1) The focus is only on how humans grasp objects (graspable affordance), rather than on which part of the tool is being used (functional affordance). (2) Affordances are learned and represented as probabilistic distributions, which are coarse and noisy, making them difficult to apply in real-life situations and susceptible to distractions.
[…]
@@
Abbildung 1: Illustration of affordance data collection from videos of human interactions. Existing work joint_hand_hotspot ; vrb collects the graspable affordance in the form of Gaussian heatmaps, whereas we extract both graspable and functional affordances with precise segmentation masks.
To resolve these limitations, we aim to jointly learn graspable and functional affordances of objects from egocentric videos, focusing on generating precise segmentation maps rather than coarse heatmaps. Specifically, to avoid costly and time-consuming manual annotation, we propose an automated pipeline to collect data without human labor. Given an egocentric video, this pipeline first extracts graspable points on objects from hand-object interactions and functional points from tool-object interactions. Since objects are often occluded by hands or tools during interaction, we identify the pre-contact frame, where contact is about to occur, using an off-the-shelf hand-object detector ho-detector . We then project the extracted points back to the pre-contact frame through homography homography or point correspondence. Finally, these graspable and functional points are employed as prompts and fed into the Segment Anything Model (SAM) sam to obtain part segmentation. We gather training data from two large-scale egocentric video datasets: Epic-kitchens epic and Ego4d ego4d .
@@
Although the data are collected in an annotation-free manner, it poses significant challenges for model training. Most training samples have quite low resolution, with cropped areas often comprising only 5% of the original frame. Additionally, factors such as motion blur, occlusion, and homography estimation can make the resulting training samples blurry and noisy. To address these issues, we propose the Geometry-guided Affordance Transformer (GAT), which includes a simple and effective Depth Feature Injector (DFI) to incorporate geometric information during training. DFI allows the model to make predictions in conjunction with 3D geometry, rather than relying solely on blurry appearances. Moreover, we observe that the model yields inferior performance when evaluated on data that significantly differs from the training source domain. To tackle this domain gap, we use the visual foundation model DINOv2 dinov2 as the image encoder, which has been trained on data from various domains. We keep DINOv2 frozen and employ Low-Rank Adaptation (LoRA) lora for fine-tuning to enhance the model’s generalization ability and prevent overfitting. After training, we combine GAT with a grasp generation model for robotic manipulation, and name this framework Aff-Grasp. Aff-Grasp can adaptively grasp objects based on different task requirements, and identify the functional affordance of objects to complete tasks.
@@
To comprehensively demonstrate the effectiveness of our data collection and model training, we perform evaluations from two perspectives. First, we collect and annotate 721 images from several existing affordance datasets and internet sources, creating a challenging evaluation dataset of great diversity to assess the model’s performance. Second, we design a real-world robotic manipulation evaluation with 7 tasks and 34 diverse objects. A task is considered successful if the model makes the correct affordance segmentation, the robot grasps the correct graspable part, and applies the functional part to the target object.
Overall, the contributions of this work can be summarized as follows:
  1. 1.
    Automated Affordance Data Collection: We propose an automated pipeline for collecting and annotating affordance data from egocentric human-object interaction videos. Different from previous work, the data are collected with precise segmentation maps for both graspable and functional affordances.
  2. 2.
    Advanced Affordance Learning Model: We introduce Geometry-guided Affordance Transformer (GAT) that features an innovative Depth Feature Injector (DFI) to incorporate geometric information. GAT builds on the visual foundation model DINOv2 and employs LoRA for fine-tuning.
  3. 3.
    Affordance-Oriented Manipulation: We design a framework, Aff-Grasp, that integrates GAT with a grasp pose generation model to achieve affordance-oriented manipulation. Given a task, Aff-Grasp can localize the most appropriate object (without specifying the object category), grasp the correct part, and utilize its functional part to complete the task.
  4. 4.
    Comprehensive Vision and Robot Evaluations: We conduct comprehensive experiments on both static datasets and real robots. A challenging affordance evaluation dataset is created for vision evaluation and affordance-oriented manipulation tasks are designed for robot experiments.
@@ 2 Related Work
2.1 Visual Affordance Learning
Affordance learning studies the properties of objects and environments that suggest possible actions. It has garnered considerable attention in the fields of computer vision and robotics. Initial successes in this field were achieved through fully supervised methods cychuang2017learning ; cad120 ; tool_parts_umd ; affordancenet ; demo2vec that employ convolutional neural networks (CNNs). These methods, however, heavily depend on large-scale, annotated datasets 3d-affordancenet ; nguyendetecting ; affordancenet ; learn2act_properly ; aff_with_CNN_iit ; bay-aff , which are costly and time-consuming to produce.
To reduce annotation costs, recent interest has shifted to weakly-supervised approaches, utilizing keypoints weakaff2 ; weakly_supervised_affordance_detection and image-level labels locate ; ag_from_exocentric_imgs ; grounded . Although these weakly-supervised based methods have made notable progress, they still require substantial manual collection of training data. Distinct from the above methods, we generate the affordance training data in an automatic manner, eliminating the need for manual annotation efforts. Also, we incorporate 3D geometry information to facilitate affordance learning, which has been shown to be beneficial in related affordance literature tool_parts_umd ; affordancellm ; rgbd-aff-video . Although geometry information like depth maps is not directly available for arbitrary images, recent monocular depth estimation models midas ; dpt ; zoedepth ; marigold ; depth-any have become much more reliable through better modeling and data-driven schemes. In this work, we use a state-of-the art depth estimation model, Depth-Anything depth-any , to produce pseudo depth maps.
2.2 Affordance Learning from Human Videos
An emerging and promising alternative for affordance learning is the automatic extraction of affordance knowledge through the observation of human interactions with objects in natural environments. Given the abundance of human-object interaction video datasets epic ; epic_v2 ; Damen2018EPICKITCHENS ; oh2019video ; ego4d ; Li_2018_ECCV ; hoi4d ; oakink ; oakink2 , recent work has explored how to extract rich affordance-based information from these videos. Liu el al. joint_hand_hotspot first proposed an automatic pipeline to generate data from egocentric videos for training, and represent affordance as probabilistic distributions in the form of interaction heatmaps. Following this paradigm, VRB vrb further collected trajectory waypoints to estimate the post-contact direction of movement. Robo-ABC robo-abc created an affordance memory consisting of object images and human contact points, achieving zero-shot generalization via object retrieval and semantic correspondence mapping.
@@ 2.2 Affordance Learning from Human Videos
Despite making significant strides in affordance learning, these studies have two limitations. First, the defined affordance is limited to graspable areas of objects, ignoring the important functional parts. Second, the affordance is predicted as coarse heatmaps, which are less accurate when applied in the real world. To address these limitations, we introduce an improved affordance data collection pipeline that localizes both graspable and functional points, and utilizes SAM sam to produce high-quality masks with point-based prompting.

### v2

# Learning Precise Affordances from Egocentric Videos for Robotic Manipulation

@@
Abstract
Affordance, defined as the potential actions that an object offers, is crucial for embodied AI agents. For example, such knowledge directs an agent to grasp a knife by the handle for cutting or by the blade for safe handover. While existing approaches have made notable progress, affordance research still faces three key challenges: data scarcity, poor generalization, and real-world deployment. Specifically, there is a lack of large-scale affordance datasets with precise segmentation maps, existing models struggle to generalize across different domains or novel object and affordance classes, and little work demonstrates deployability in real-world scenarios. In this work, we address these issues by proposing a complete affordance learning system that (1) takes in egocentric videos and outputs precise affordance annotations without human labeling, (2) leverages geometric information and vision foundation models to improve generalization, and (3) introduces a framework that facilitates affordance-oriented robotic manipulation such as tool grasping and robot-to-human tool handover. Experimental results show that our model surpasses the state-of-the-art by 13.8% in mIoU, and the framework achieves 77.1% successful grasping among 179 trials, including evaluations on seen, unseen classes, and cluttered scenes. Project page: https://reagan1311.github.io/affgrasp.
@@ 1 Introduction
Figure 1: Illustration of affordance data collection and robot deployment. Existing work [27, 28, 19, 2] collects the graspable affordance as Gaussian heatmaps, whereas we extract both graspable and functional affordances with precise segmentation masks, enabling tool grasping, tool-object interaction, and robot-to-human tool handover.
Understanding affordances of an object means knowing the possible actions it enables as well as locating specific object parts for those actions. This knowledge serves as a key link between perception and action, enabling an intelligent system to transition from observing how an object is used to actually performing the action itself. Despite its importance and recent progress, current affordance-related research often suffers from three major challenges: data scarcity, poor generalization, and real-world deployment. First, affordance datasets are scarce and expensive to create, as they require detailed annotations of object parts, which are small, low-resolution, and often occluded. For example, accurately segmenting the handle of a spoon for holding is challenging due to its thin structure and potential occlusions in cluttered scenes. Second, current models often struggle to generalize to diverse new objects, affordances, or environments, as they are typically trained from scratch on a limited set of data. Finally, few affordance studies deploy methods in real robots, where the models need to be robust to noise, novel scenes, and other real-world factors. These three challenges are in fact deeply interconnected: a lack of large-scale and diverse data limits model generalization, which is essential for reliable real-world deployment.
To tackle each of these challenges, we propose a comprehensive affordance learning system that encompasses data collection, model learning, and robot deployment. The system starts with an automatic pipeline for collecting training data from egocentric videos, which are rich in human-object interactions. While a number of studies [36, 28, 2, 19] have explored extracting affordances from egocentric videos, two limitations persist in their pipeline as illustrated in Fig. 1: (1) The focus is primarily on how humans grasp objects (graspable affordance), rather than on which part of the tool is being used (functional affordance). (2) Affordances are learned and represented as probabilistic distributions, which are coarse and noisy, making them difficult to apply in real-life situations and susceptible to distractions. To resolve these limitations, we aim to jointly annotate both graspable (e.g., object handles) and functional affordances (e.g., knife blades, hammerheads), focusing on generating precise segmentation maps rather than coarse heatmaps. Concretely, given an egocentric video, our pipeline first extracts graspable points on objects from hand-object interactions and functional points from tool-object interactions. To cope with occlusions, we identify the pre-contact frame, where contact is about to occur, and project the extracted points to this frame through homography or point correspondence. Finally, these points act as prompts for the segment anything model (SAM) [21] to obtain precise part segmentation.
@@
Although the previous step enables collection of images and annotations at a low cost, the resulting data pose challenges for training. Most samples have quite low resolution, with cropped areas often comprising only 5% of the original frame. To address this issue, we propose Geometry-guided Affordance Transformer (GAT), which enriches and constrains the prediction process by leveraging the geometric features, rather than relying solely on blurry or low-resolution appearance. Moreover, we observe that the model yields inferior performance when evaluated on data that largely differs from the training source domain. To tackle this domain gap, we use the visual foundation model DINOv2 [42] as the image encoder, which has been trained on data from various domains. Lastly, to enable affordance-oriented manipulation, we introduce Aff-Grasp, a framework that combines the GAT with a grasp generation model for robotic manipulation. Building on the accurate predictions of graspable and functional affordances by GAT, Aff-Grasp can handle tasks beyond simple grasping, including tool-object interaction and robot-to-human tool handover (see examples in Fig. 1).
@@
To comprehensively demonstrate the effectiveness of our methods, we perform evaluations from two perspectives. First, we collect and annotate images from several existing affordance datasets and internet sources, creating a challenging evaluation dataset of great diversity to assess the model’s performance. Second, we design a real-world robotic manipulation evaluation with 7 tasks and 34 diverse objects. It is worth nothing that both evaluations include out-of-domain data and novel objects, making this a challenging cross-domain and zero-shot setup.
Overall, the contributions of this work can be summarized as follows: (1) Automated Affordance Data Collection: We propose an automated pipeline for collecting and annotating affordance data from egocentric human-object interaction videos. Different from previous work, the data are collected with precise segmentation maps for both graspable and functional affordances. (2) Advanced Affordance Learning Model: We introduce Geometry-guided Affordance Transformer (GAT) that incorporates shape and geometric priors in a flexible and innovative way to tackle the challenging affordance segmentation task. (3) Affordance-Oriented Manipulation: We present Aff-Grasp, a framework designed for affordance-oriented grasping. Aff-Grasp can identify the most suitable object based on task instructions, grasp the appropriate part, and utilize its functional component to complete the task (without specifying explicit object or part names). (4) Extensive Vision and Robot Evaluations: We conduct experiments on both visual data and a real robot. A challenging affordance evaluation dataset is created for vision evaluation and diverse tasks are designed for robot experiments across a wide range of objects.
@@ 2 Related Work
Figure 2: Illustration of the data collection process from egocentric videos. First, graspable points (depicted in purple) are localized from clips of hand-object interaction and then projected to pre-contact frame by homography. Next, functional points (depicted in green) are identified from tool-object interactions and mapped to the pre-contact frame of hand-object interaction through point correspondence. Lastly, these points are used as prompts for the SAM to obtain affordance masks.
@@
Learning Affordance from Human Videos. An emerging and promising alternative for fully supervised affordance learning is the automatic extraction of affordance knowledge through the observation of human interactions. Given the abundance of human-object interaction video datasets [8, 9, 7, 39, 15, 25, 30, 56, 60], many studies [14, 27, 28, 2, 19] have explored how to extract rich affordance information from videos. Despite making significant strides in affordance learning, these studies have two limitations. First, the defined affordance is limited to graspable areas of objects, ignoring the important functional parts. Second, the affordance is predicted as coarse heatmaps, which are less accurate when applied in the real world. While some recent work [34, 59, 58] has made attempts to address these limitations, none have tackled both aspects simultaneously. To bridge these gaps, we introduce an improved affordance data collection pipeline that localizes both graspable and functional points, and utilizes SAM [21] to produce high-quality masks with point-based prompting.
Affordance-Oriented Robotic Manipulation. Recent work [18, 44, 53, 48, 52, 17, 33] has achieved task-oriented grasping using vision-language models (VLMs) and large language models (LLMs) [43, 42, 3, 46, 41, 40]. These methods use LLMs to infer parts to be grasped based on task instructions, and VLMs to localize the relevant regions. While this pipeline can produce effective and accurate manipulation, the reasoning process often requires prompt engineering and can be time-consuming. In contrast, there has been less focus on more efficient affordance-oriented grasping that can derive graspable and functional areas without explicitly specifying corresponding object parts. One major barrier to this direction is the lack of large-scale affordance datasets, compounded by the difficulty of unifying existing ones due to inconsistent annotation standards.
To overcome the data limitations, recent methods such as VRB [2] and Robo-ABC [19] have explored the potential of learning affordances from human videos for robotic manipulation. However, VRB generates only coarse Gaussian heatmaps, and Robo-ABC relies on point correspondences that can be noisy and susceptible to background variations. In contrast, we propose a more effective and robust strategy to extract precise affordance masks, and introduce an efficient model for affordance learning. When combined with a small set of basic pre-recorded motion primitives (as in [11, 10]) and grasp pose detection models [49, 13, 12], our method enables the utilization of both graspable and functional affordances.
3 Method
In this work, our goal is to develop a holistic system that covers data collection, model learning, and robot deployment. To this end, we first develop an automated pipeline to collect images and related affordance annotations from human videos. Next, we propose an effective affordance learning model termed Geometry-guided Affordance Transformer (GAT). Finally, we introduce Aff-Grasp that couples the trained model with an off-the-shelf grasp estimation model to achieve affordance-oriented manipulation. In Sec. 3.1, we describe how affordance data are collected from large-scale egocentric videos of human interactions. We then elaborate on the design of GAT that enables effective affordance learning from collected data in Sec. 3.2. Lastly, in Sec. 3.3, we explain Aff-Grasp, detailing how it yields affordance-oriented grasp poses.
Figure 3: The architecture of GAT. It consists of a DINOv2 image encoder, a depth feature injector, an embedder, and LoRA layers. The model performs segmentation by computing cosine similarity between upsampled features and learnable / CLIP text embeddings.
3.1 Data Collection from Egocentric Videos
Given an egocentric video of a human interacting with an object, our aim is to first locate contact points. Human-object interaction videos can generally be categorized into two types: hand-object interaction and tool-object interaction. In hand-object interaction, contact points indicate where the human grasps the object. In tool-object interaction, contact points reveal which part of the tool is used to interact with the target object. These points represent sparse graspable and functional areas of an object, carrying rich affordance information. As shown in Fig. 2, we propose a pipeline to automatically collect these points without manual annotation. The collected points then serve as prompts to produce precise segmentation masks using SAM [21].
Graspable Point Localization. Egocentric videos like Epic-Kitchens [8] and Ego4D [15] include timestamped narrations that describe actions and their respective start and end times. Based on these narrations, we first retrieve hand-object interaction clips (associated with actions such as “take” or “hold”) and employ a hand-object detector [47] to generate contact states and hand-object bounding boxes for all frames. Next, we use labeled timestamps to extract the contact frame, which is typically annotated as the start of an action. We conduct an additional hand segmentation in this frame using the hand box as a prompt via EfficientSAM [55]. We then locate the intersection region of the hand mask and object bounding box to sample $n$ contact points $P=\{p_{1},p_{2},...,p_{n}\}$. However, the sampled points are often occluded by hands, and therefore do not accurately represent the graspable affordance area of the object. To collect clean object images free of occlusion, it is necessary to identify the pre-contact frame, i.e., the last frame where the object is fully visible before contact occurs. We utilize the contact states to detect the frame that is closest to the contact frame but without hand-object contact, designating this as the pre-contact frame. Since human motion between adjacent frames is minimal, we follow a similar pipeline to previous studies [28, 19] to project the average position of sampled graspable points to the pre-contact frame by computing a homography transformation [50], as illustrated by the purple dashed line in Fig. 2.
Functional Point Localization. To localize functional points, we first retrieve the relationship between objects and affordances from existing affordance datasets. We then extract related tool-object interaction clips from video narrations based on the object-affordance relationship. For example, most datasets associate affordances “cut” and “grasp” with a knife. After localizing the graspable points from a clip showing a person grasping a knife, we then retrieve the following nearest clip depicting a cutting action with the knife.
However, the tool is often heavily occluded or invisible in the contact frame. Similar to graspable point localization, we need to find the pre-contact frame that shows minimal or no intersection between the tool and the target object. To achieve this, we first employ the same combination of the hand-object detector and EfficientSAM to obtain the bounding box and mask of the hand-held tool. Next, we use an open-vocabulary object segmentation model, GroundedSAM [45], to segment the target object. We then measure the Intersection over Union (IoU) between tool and object bounding boxes in frames prior to the contact frame until the IoU is below a preset threshold. Lastly, we calculate the point distances between masks in this pre-contact frame and extract the point within the tool mask that has the shortest distance to all points in the object mask. An erosion operation is applied to the tool mask to ensure that the functional point is inside the tool.
Nonetheless, not all object categories have related action clips in the narrations. In such cases, we use the farthest sampling to determine the functional points based on the distance to the grasp points. This simple method produces accurate functional points, as most tools are designed with graspable and functional parts distributed at opposite ends.
Data Generation. After extracting functional points, we first project these points to the pre-contact frame of the hand-object interaction clip where we infer the graspable points. Since the object category remains the same, we compute the point correspondence within object bounding boxes using foundation model features [1], which map the functional point from the tool-object pre-contact frame to the hand-object pre-contact frame (illustrated by the green dashed line in Fig. 2). We then label the graspable points as positive and the functional points as negative to obtain the graspable affordance mask. Conversely, the functional affordance mask is generated by reversing the positive and negative labels. Finally, we crop the object images and store them along with the generated segmentation masks as annotations.
3.2 Geometry-guided Affordance Transformer
While the affordance data are collected from egocentric videos without manual labor, there are two issues that hinder effective model training. The first issue is the low resolution of the collected images. The object of interest occupies very small areas of the video frames, often resulting in cropped images of less than 100 pixels in either height or width. The second is the limited diversity of the training data, characterized by monotonous backgrounds and mostly restricted to indoor scenes. To cope with these issues, we propose an affordance learning architecture called GAT, as illustrated in Fig. 3. It includes a novel Depth Feature Injector (DFI) that integrates geometric priors into image features using pseudo depth maps, a DINOv2 image encoder as a feature extractor, and additional LoRA layers for effective fine-tuning.
Depth Feature Injector. We argue that depth maps introduce rich geometric information that can help with foreground-background and part separation. Additionally, they exclude color information, allowing the model to fully focus on the shape information, which is highly relevant to affordances. For instance, graspable parts often consist of shapes like cylinders or spheres, while parts designed for cutting typically feature sharp edges and flat surfaces.
Specifically, we first obtain pseudo depth maps for each training image with a state-of-the-art depth estimation model Depth-Anything [57]. During model training, the pseudo depth map is first encoded into feature maps using a stem block, which contains three standard $3\times 3$ convolution layers. These feature maps are then processed by a $1\times 1$ convolution that transforms the channel dimension to match that of the RGB image features.
$Q=\phi_{q}(F_{i}),~K=\phi_{k}(F_{d}),~V=\phi_{v}(F_{d}),$ (1)
$\hat{F}_{i}=\beta\cdot\text{softmax}(QK^{T}/\sqrt{d_{k}})\cdot V+F_{i},$ (2)
where $\phi$ is a linear transformation, and $d_{k}$ is the dimension of the key acting as a scaling factor. Following [5], we set a learnable vector $\beta\in\mathbb{R}^{C}$, initialized to 0, to balance the output from the cross-attention layer and the image feature. This strategy prevents the image feature from being excessively affected by the depth feature, making the training process more stable. We observe that DFI constantly brings improvement, even when integrated solely during training (see Sec. 4.3). This indicates that it can act as a regularization mechanism during training, and can be discarded during inference to speed up the process.
DINOv2 with Low-Rank Adaptation. We notice that directly training from the typical ImageNet pre-trained representation often leads to inferior results. This can be attributed to two primary reasons: First, affordance segmentation focuses on fine-grained object parts, whereas representations trained for image classification emphasize global object features [1]. Second, ImageNet pre-trained models exhibit limited diversity, making it challenging to handle data from diverse domains. To address this issue, we employ the self-supervised visual foundation model DINOv2, which has been demonstrated to be highly effective for data-limited affordance learning [24]. Furthermore, we introduce LoRA [16] to fine-tune the model without modifying the parameters of the original DINOv2. This strategy helps adaptation across different domains and prevents overfitting. LoRA was originally developed to fine-tune large language models for different downstream tasks. Specifically, it injects trainable rank decomposition matrices to a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ by $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During training, only $A$ and $B$ are trainable, while $W_{0}$ remains frozen. This incurs minimal computational cost and memory usage.
Figure 4: The framework of Aff-Grasp. It first employs an open-vocabulary detector [29] to locate all objects within the scene, which are then sent to GAT to determine if they possess the corresponding affordance required for the task. Afterwards, a 6 DoF grasp generation model, Contact-GraspNet, leverages the object’s graspable affordance and the depth map to generate dense grasp proposals. Finally, the robot executes affordance-specific sequential motion primitives to apply the functional part to the target.
Classifier and Loss Functions. To speed up inference time for real-world applications, we avoid adding any complex decoder structures. Instead, we process the output feature with an embedder (an MLP), reshape it, and upsample it by a factor of four to increase the resolution $F_{out}\in\mathbb{R}^{C\times\frac{4H}{p}\times\frac{4W}{p}}$, where $p$ is the patch size. Next, we intialize $M\in\mathbb{R}^{{L\times C}}$ learnable embeddings, where $L$ is the number of affordance categories. We compute the cosine similarity between $M$ and $F_{out}$ to yield the segmentation output, which is then restored to the same size as the input image via bilinear interpolation. Due to the domain gap, we do not add a learnable embedding for the background classification to prevent overfitting. Alternatively, we determine a pixel as background if all its affordance predictions are below a preset threshold $\tau$. Compared to a linear layer and explicit background classifier, the cosine similarity-based segmentation and implicit background prediction are more robust and can effectively improve the performance, as detailed in Sec. 4.3. In addition, to achieve open vocabulary affordance segmentation, $M$ can also be replaced with corresponding CLIP text embeddings, as verified in [24].
Since the collected data are highly unbalanced, we utilize a combination of focal loss [26] and dice loss [32] as training objectives:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{focal}+\mathcal{L}_{dice},$ (3)
where $\alpha$ is a weighting factor to balance loss values.